The product of the ages of A and B is 104. If thrice the age of B is less than A's age by 11 years. Find the B's age.thanks
Answer by checkley77(12844)

AB=104
3B=A-11 OR A=3B+11
(3B-11)B=104
3B^2-11B-104=0
(3B+13)(B-8)=0
B-8=0
B=8 AGE OF B.
A*8104
A=104/8=13 AGO OF A.
PROOF:
13=3*8-11
13=23=11
13=13